# Classification of compact manifolds with positive isotropic curvature

### v6

@@ 3 Existence of $(r,\delta)$-surgical solutions and applications
Proof of Corollary 1.3. Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature and with $|\pi_{1}(M)|<\infty$. By the assumptions on $M$, we see that there is at most one nontrivial summand in the connected sum decomposition of $M$ given by the conclusion of Theorem 1.1, which must be diffeomorphic to a spherical space form. $\Box$
@@
Proof of Corollary 1.5. Let $M$ satisfy the assumption of Corollary 1.5. By Corollary 1.2 in [19], which may also be viewed as a corollary of Theorem 1.1 here, there is a finite cover of $M$ which is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. From [28] we have an explicit construction of the universal cover, denoted by $\widetilde{M}$, of a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Observe that the image of a sphere $\mathbb{S}^{k}$ under a continuous map from $\mathbb{S}^{k}$ to $\widetilde{M}$ is compact for any $k\in\mathbb{N}$. By using induction, the Blakers-Massey theorem, and the long exact sequence for the relative homotopy groups of a pair, we see that $\pi_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2. Alternatively, once we know the universal cover $\widetilde{M}$, we can compute the homology groups $H_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2 (by using the Mayer-Vietoris sequence), and conclude $\pi_{i}(\widetilde{M})=0$ for $22\leq i\leq n-$ 2 by using the Hurewicz theorem. Now the corollary follows. $\Box$
@@
Proof of Corollary 1.6. Let $M$ satisfy the assumption of Corollary 1.6. From the proof of Corollary 1.2 in [19], $M$ admits a finite sheeted regular cover, denoted by $\tilde{M}$, which falls into two cases: (1) $\tilde{M}$ is orientable, and is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$; (2) $\tilde{M}$ is non-orientable, and is diffeomorphic to $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ and $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. In the first case we have $H_{i}(\tilde{M})=0$ for $22\leq i\leq n-$ 2 by using the Mayer-Vietoris sequence, and
@@
(cf. for example, Example 2.6 in [15],) where $G$ is the deck transformation group of the covering map $\tilde{M}\rightarrow M$. In the second case, we take the orientation double cover, denoted by $\tilde{\tilde{M}}$, of $\tilde{M}$. Of course, an orientable double cover is a regular cover. Moreover we know that $\tilde{\tilde{M}}$ is diffeomorphic to $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Now we can argue as in case (1) to conclude $H^{i}(\tilde{M};\mathbb{R})=0$ and $H^{i}(M;\mathbb{R})=0$ for $22\leq i\leq n-$ 2. $\Box$
Remark. In a preprint posted on the arXiv in January 2008, S. Gadgil and H. Seshadri pointed out that if the fundamental group of a compact manifold $M$ of dimension $n\geq 4$ with positive isotropic curvature is virtually free, then $b_{i}(M)=0$ for $22\leq i\leq n-$ 2.

### v8

@@ 3 Existence of $(r,\delta)$-surgical solutions and applications
Proof of Corollary 1.3. Let $(M,g_{0})$ be a compact connected manifold of dimension $n\geq 12$ with positive isotropic curvature and with $|\pi_{1}(M)|<\infty$. By the Seifert-van Kampen theorem, we see that there is at most one nontrivial summand in the connected sum decomposition of $M$ given by the conclusion of Theorem 1.1, which must be diffeomorphic to a spherical space form. $\Box$
@@
Proof of Corollary 1.4. Let $M$ satisfy the assumption of Corollary 1.4. By Corollary 1.2 in [27], which may also be viewed as a corollary of Theorem 1.1 here, there is a finite cover of $M$ which is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. From [38] we have an explicit construction of the universal cover, denoted by $\widetilde{M}$, of a connected sum of a finite number of copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Observe that the image of a sphere $\mathbb{S}^{k}$ under a continuous map from $\mathbb{S}^{k}$ to $\widetilde{M}$ is compact for any $k\in\mathbb{N}$. By using induction, the Blakers-Massey theorem, and the long exact sequence for the relative homotopy groups of a pair, we see that $\pi_{i}(\widetilde{M})=0$ for $2\leq i\leq n-2$. Alternatively, once we know the universal cover $\widetilde{M}$, we can compute the homology groups $H_{i}(\widetilde{M})=0$ for $2\leq i\leq n-2$ (by using the Mayer-Vietoris sequence), and conclude $\pi_{i}(\widetilde{M})=0$ for $2\leq i\leq n-2$ by using the Hurewicz theorem. Now the corollary follows. $\Box$
@@
Proof of Corollary 1.5. Let $M$ satisfy the assumption of Corollary 1.5. From the proof of Corollary 1.2 in [27], $M$ admits a finite sheeted regular cover, denoted by $\tilde{M}$, which falls into two cases: (1) $\tilde{M}$ is orientable, and is diffeomorphic to $\mathbb{S}^{n}$, or $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$; (2) $\tilde{M}$ is non-orientable, and is diffeomorphic to $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\tilde{\times}\mathbb{S}^{1}$ and $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. In the first case we have $H_{i}(\tilde{M})=0$ for $2\leq i\leq n-2$ by using the Mayer-Vietoris sequence, and
@@
(cf. for example, Example 2.6 in [22],) where $G$ is the deck transformation group of the covering map $\tilde{M}\rightarrow M$. In the second case, we take the orientation double cover, denoted by $\tilde{\tilde{M}}$, of $\tilde{M}$. Of course, an orientable double cover is a regular cover. Moreover we know that $\tilde{\tilde{M}}$ is diffeomorphic to $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$, or a connected sum of finitely many copies of $\mathbb{S}^{n-1}\times\mathbb{S}^{1}$. Now we can argue as in case (1) to conclude $H^{i}(\tilde{M};\mathbb{R})=0$ and $H^{i}(M;\mathbb{R})=0$ for $2\leq i\leq n-2$. $\Box$
Remark. In a preprint posted on the arXiv in January 2008, S. Gadgil and H. Seshadri pointed out that if the fundamental group of a compact manifold $M$ of dimension $n\geq 4$ with positive isotropic curvature is virtually free, then $b_{i}(M)=0$ for $2\leq i\leq n-2$.